the second angle of a triangular kite is four times as large as the first. The third angle is 5 degrees more than the sum of the other two angles. Find the measure of the second angle.
Answer by mananth(16946)

let first angle be x
the second angle = 4x
..
Third angle =(x+4x)+5
....
x+4x+5x+5 =180
10x+5 =180
-5
10x+5-5 = 180-5
10x = 175
/10
x= 175/10
x =17.5 first angle
second angle = 4x = 4*17.5 = 70 deg
third angle = 5x+5 = 5*17.5 + 5= 92.5 deg